A box of weight X is allowed to slide an incline of a°. If the coefficient of kinetic friction between block and incline is 0.1


Find the acceleration down the incline


1
Expert's answer
2021-03-31T17:06:10-0400

"m\vec{a}=m\vec{g}+\vec{N}+\vec{F},"

"ma=mgsin\alpha-\mu mg cos\alpha,"

"a=gsin\alpha-0.1 g \cos \alpha."